OpenStudy (anonymous):

find the solution set by factoring and using zero product rule or by using the quadratic formula 12x^2-12x=1

OpenStudy (anonymous):

shud try solving urself

OpenStudy (anonymous):

i've tried itz confusing

OpenStudy (anonymous):

anyway \[ x = 1/6 (3+2 \sqrt(3))\] \[x = 1/6 (3-2 \sqrt(3))\]

OpenStudy (anonymous):

is it okay